# Elucidated Rolling Diffusion Models for Probabilistic Weather Forecasting

### v1

@@ Results.
An evaluation of ERDM’s performance, illustrated in Fig. 5, reveals its consistent superiority in relative CRPS skill over our primary EDM baseline, by up to $10\%$. Similarly, it significantly outperforms Graph-EFM on the four CRPS scores reported in their study [37]. Compared to the more computationally complex models, IFS ENS and NeuralGCM, ERDM performs competitively, but exhibits short-range weakness on some variables versus IFS ENS. We attribute this to our EDM-based initialization strategy, which could potentially be addressed by initializing the first window of ERDM with IFS ENS. In terms of spread-skill ratio, ERDM, along with IFS ENS, delivers the most calibrated forecasts (see Fig. 12). Conversely, EDM and NeuralGCM often show underdispersion in short-range forecasts. A crucial aspect of evaluating any ensemble forecast is its physical realism. In this regard, ERDM performs exceptionally well. As depicted in Fig. 6, its normalized power spectra are on par with those of the physics-based IFS ENS. This is a significant achievement, as many ML models typically struggle to reproduce such physically consistent spectra—a challenge evident with NeuralGCM in Fig. 6 and documented by Rasp et al. [43]. Results for more variables, absolute power spectra, spread-skill ratio results, as well as forecast visualizations, are included in Appendix F.
@@ 5.4 Ablations
We conducted a comprehensive series of ablation studies on the Navier-Stokes forecasting task to validate our key design choices and understand their individual contributions to ERDM’s performance. We analyze it in full detail in Appendix E, and only explain the crucial design choices here. Our ablations reveal that certain architectural and training strategy elements are fundamental to ERDM’s success, with their absence or significant alteration leading to performance substantially worse than even the single-step EDM baseline.
@@
    Appropriate progressive noise schedule & fixed training: The design of the noise schedule is paramount. Default EDM hyperparameters (e.g., $\rho=7$) prove suboptimal for ERDM, yielding 2×2$\times$2 × worse CRPS than our proposed default schedule with $\rho=-10$. See Fig. 2 for intuition on why. Furthermore, training with a fixed, optimized noise schedule is critical; randomizing the schedule during training degrades performance by nearly 2×2$\times$2 ×. Despite this, such randomized training can serve as a heuristic to identify an effective fixed schedule for subsequent retraining. This finding contrasts with the randomized training procedure from Chen et al. [9].
@@
    Strategic loss weighting: Reweighing the losses based on the lognormal probability density function of noise levels, $f(\sigma)$, as proposed, is crucial. Removing this weighting causes a >2×>2$\times$> 2 × performance drop. We generally found that a larger mean of the lognormal distribution, $P_{\text{mean}}>0$, than EDM’s default, $-1.2$, was necessary for ERDM to achieve optimal results, confirming the importance of tuning here.
@@
    Dedicated spatiotemporal architecture: A crucial element is the use of a proper architecture that explicitly models spatiotemporal dependencies. Naively stacking the time dimension into the channel dimension of a 2D architecture results in severe 4×4$\times$4 × performance degradation, confirming the need for our bespoke temporal architecture.
@@ 6 Conclusion
In this work, we introduced Elucidated Rolling Diffusion Models (ERDM), a novel framework designed to address the significant challenges of long-range forecasting in complex, chaotic scientific systems. By adapting and extending the successful design of EDM, ERDM incorporates several key innovations: a progressive temporal noise schedule to explicitly model increasing forecast uncertainty, snapshot-dependent preconditioning and loss weighting tailored for sequential data, an efficient decoupled initialization strategy, and a bespoke hybrid spatiotemporal architecture. Comprehensive ablation studies confirm the need for synergistically combining these design choices.
Our extensive experiments on the demanding ERA5 weather dataset and Navier-Stokes flow simulations demonstrate ERDM’s efficacy at probabilistic forecasting. It consistently outperforms relevant EDM-based baselines, particularly in terms of CRPS and calibration over extended forecast horizons. Although ERDM achieves these results with fewer neural function evaluations (e.g., $12$ sampling steps per snapshot on ERA5 vs. $20$ for the EDM baseline), the greater computational cost per evaluation due to its 3D architecture results in sampling times similar to next-step EDM forecasting.
ERDM offers a robust and reliable approach for probabilistic forecasting in scientific domains. Future work could focus on optimizing the computational efficiency of the 3D architecture, exploring dynamic adjustments of the window size and synergies with latent space modeling, and applying the ERDM framework to a broader range of challenging spatiotemporal dynamics.
@@ Acknowledgements
S.R.C. acknowledges generous support from the summer internship and collaboration at NVIDIA. This research used resources from the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility, using NERSC awards DDR-ERCAP0034142, ASCR-ERCAP0033209, and EESSD-ERCAP0033799. This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, and NSF Grants #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069, DARPA AIE FoundSci and DARPA YFA.
@@ Appendix B Background on Diffusion Models
Conversely, the backward diffusion process operates by initially sampling noise, represented as ${{\bm{x}}}_{0}$, from a prior distribution, typically $\mathcal{N}(\mathbf{0},\mathbf{G}(0))$, where $\mathbf{G}(0)$ corresponds to the maximum noise level. The process then focuses on denoising this sample through a sequence, ${{\bm{x}}}_{i}$, characterized by a sequence of decreasing noise covariance matrices: $\mathbf{G}(t_{0})\succ\mathbf{G}(t_{1})\succ\ldots\succ\mathbf{G}(t_{N})\approx\mathbf{0}$ (in the Loewner order), where $t_{0}=0$. Each intermediate sample ${{\bm{x}}}_{i}$ is drawn from ${{\bm{x}}}_{i}\sim p({{\bm{x}}}_{i};\mathbf{G}(t_{i}))$. The terminal sample of the backward process, ${{\bm{x}}}_{N}$, is expected to approximate the original data distribution $p_{{\rm data}}({{\bm{y}}})$. Note that for the sake of consistency with our ERDM formulation, we reversed the diffusion time order. That is, $t=0$ ($1$) corresponds to maximal (minimal) noise levels, which is sometimes described in opposite order in the diffusion literature.
SDE formulation. To present the forward and backward processes rigorously, they can be captured via stochastic differential equations (SDEs). Such SDEs ensure that the sample, ${\bm{x}}(t)$, aligns with the designated data distribution, $p({\bm{x}};\mathbf{G}(t))$, as it evolves through diffusion time $t$ [51]. The Elucidated Diffusion Model (EDM) by Karras et al. [23] provides a principled design for diffusion models, typically based on a scalar noise schedule $\sigma(t)$ (leading to isotropic noise $\sigma^{2}(t)\mathbf{I}$). The formulation presented here extends these principles to accommodate a general, potentially anisotropic, time-varying noise covariance matrix $\mathbf{G}(t)$. Let $\dot{\mathbf{G}}(t)=d\mathbf{G}(t)/dt$ be the time derivative of the noise covariance matrix. We assume $\dot{\mathbf{G}}(t)$ is positive semi-definite for all $t$, and let $\mathbf{K}(t)$ be any matrix such that $\mathbf{K}(t)\mathbf{K}(t)^{T}=\dot{\mathbf{G}}(t)$ (e.g., via Cholesky decomposition if $\dot{\mathbf{G}}(t)$ is strictly positive definite, or other matrix square roots). The forward SDE, describing the process of adding noise such that ${\bm{x}}(t)-{\bm{x}}(1)\sim\mathcal{N}(\mathbf{0},\mathbf{G}(t))$ (assuming $\mathbf{G}(1)=\mathbf{0},$ and ${\bm{x}}(1)\sim p_{\text{data}}({\bm{y}})$), is:
@@
where $\boldsymbol{\omega}(t)$ is a standard $d$-dimensional Wiener process. The corresponding backward (reverse-time) SDE is given by [51]:
@@
where $\Delta\mathbf{G}(\tau_{i})=\mathbf{G}(\tau_{i})-\mathbf{G}(\tau_{i+1})>\mathbf{0}$, we have used the approximation $\dot{\mathbf{G}}(\tau_{i})\Delta\tau_{i}\approx\mathbf{G}(\tau_{i})-\mathbf{G}(\tau_{i+1})$ (assuming $\Delta\mathbf{G}(\tau_{i})$ is small), and $\mathbf{z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a standard Gaussian noise sample. The term ^$\hat{\nabla}_{{{\bm{x}}}}\log p({\bm{x}}_{\tau_{i}};\mathbf{G}(\tau_{i}))$ is the score estimated using the trained denoiser:
^$\hat{\nabla}_{{{\bm{x}}}}\log p({\bm{x}}_{\tau_{i}};\mathbf{G}(\tau_{i}))=(\mathbf{G}(\tau_{i}))^{-1}(D_{\theta}({\bm{x}}_{\tau_{i}};\mathbf{G}(\tau_{i}))-{\bm{x}}_{\tau_{i}}).$ (10)
More sophisticated samplers, such as those proposed by Karras et al. [23] (e.g., second-order Heun or DPM-Solver++), can also be adapted for the general covariance case to improve sample quality or reduce the number of sampling steps $N$. These often involve more complex update rules, potentially incorporating predictor-corrector steps or higher-order approximations of the SDE.
@@ C.1 Probability flow ODE derivation
Recall that ERDM operated on noisy sequences $\bar{\bm{x}}_{1:W}$ of size $W$, with each $\bar{\bm{x}}_{w}\in\mathbb{R}^{D}$ corrupted according to increasing noise level standard deviations $0\approx\bar{\sigma}_{1}(t)<\bar{\sigma}_{2}(t)<\dots<\bar{\sigma}_{W}(t)$. The noise covariance matrix for the entire window $\bar{\bm{x}}_{1:W}$ at diffusion time ${t}$ is block-diagonal:
@@
where $\boldsymbol{\omega}({t})$ is a standard $WD$-dimensional Wiener process. The corresponding backward (reverse-time) SDE, which is the primary SDE used for generation, is [51]:
@@
where $D_{\theta}(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))_{w}$ is the $w$-th snapshot output of the denoiser $D_{\theta}$, which is conditioned on the full noisy window $\bar{\bm{x}}$ and the vector of current noise standard deviations $\bm{\bar{\sigma}}({t})$. The probability flow ODE, which provides a deterministic path for generation, is obtained by removing the stochastic term from the backward SDE and adjusting the drift (see Song et al. [51], Eq. (14) and Karras et al. [23], Eq. (5)):
@@
Using the specific EDM parameterization where the ODE drift is $-\mathbf{D}({t})\nabla_{\bar{\bm{x}}}\log p(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))$ with $\mathbf{D}({t})=\text{diag}(\bar{\sigma}_{1}({t})\dot{\sigma}_{1}({t})\mathbf{I}_{D},\ldots,\bar{\sigma}_{W}({t})\dot{\sigma}_{W}({t})\mathbf{I}_{D})$, the ODE becomes:
@@ C.2 Sampling
ERDM generates sequences with a sliding window of size $W$. Generating each window requires solving the backward SDE in equation 17 (or ODE in equation 20). In each iteration $k$ where the SDE is fully solved, ERDM operates on the noisy window $\bar{\bm{x}}^{(k)}_{1:W}$. After solving the SDE backwards $k\geq 1$ times, the first snapshot of the window, $\bar{\bm{x}}^{(k)}_{1}$ (with relative window index $w=1$), corresponds to a forecast of ${\bm{y}}_{k}$, where $k$ is the absolute time step of the forecast time dimension. Solving the SDE requires discretizing it into $N_{k}$ diffusion “time” steps, each of size $\Delta{t}_{k}:=\frac{1}{N_{k}}$. In general, $N_{k}$ may vary across iterations. To simplify our illustrations of the sampling process in the following, we assume that $N_{k}$ and $\Delta{t}_{k}$ are identical for all $k$, denoted as $N$ and $\Delta{t}$ respectively. This can be achieved by setting $N$ in our sampling algorithms to a positive integer. Note that in practice $N$ need not be an integer (e.g., we use $N=1.25$ for our Navier-Stokes experiments), which is accounted for in our sampling algorithms. With this simplification, the SDE evolves, for all $k$, following $\bm{\bar{\sigma}}(t_{1})\rightarrow\bm{\bar{\sigma}}(t_{2})\rightarrow\dots\rightarrow\bm{\bar{\sigma}}(t_{N})$, where $t_{i+1}=t_{i}+\Delta{t}$ for $i\geq 1$, $t_{1}=0$, and $t_{N}=1$. For example, if $N=1$ the SDE is solved from $\bm{\bar{\sigma}}(0)$ to $\bm{\bar{\sigma}}(1)$ in one sampling step, while if $N=2$ it would be solved following $\bm{\bar{\sigma}}(0)\rightarrow\bm{\bar{\sigma}}(\frac{1}{2})\rightarrow\bm{\bar{\sigma}}(1)$ in two sampling steps (see illustration in Fig. 1). At the end of the $N$ sampling steps, $\bar{\sigma}_{1}(1)=\sigma_{\text{min}}$ and the first snapshot in the window can be emitted as the forecast for lead time $k$, $\bar{\bm{x}}^{(k)}_{1}\approx{\bm{y}}_{k}$. In the next iteration of $N$ sampling steps, the next snapshot $k+1$ will be predicted by operating on a new window, $\bar{\bm{x}}^{(k+1)}_{1:W}$. Because $\bar{\sigma}_{w+1}(1)=\bar{\sigma}_{w}(0)$ for $w\geq 1$, this new window can be constructed by shifting the noisy parts of the old window by one position. That is, $\bar{\bm{x}}^{(k+1)}_{w}:=\bar{\bm{x}}^{(k)}_{w+1}$ for $w<W$, and a new pure-noise snapshot is appended as $\bar{\bm{x}}^{(k+1)}_{W}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I})$. The new window is thus noised according to standard deviations $\bm{\bar{\sigma}}(0)$ and the process can be restarted and repeated indefinitely, as illustrated in Fig. 1.
@@
The ODE is thus solved in parallel for all snapshots in the window. When $\bar{\bm{x}}(N)$ is reached, its noise levels correspond to $\bm{\bar{\sigma}}(N\Delta{t})=\bm{\bar{\sigma}}(1)$. Thus, the first snapshot is emitted, and the window is shifted as describe above. Higher-order stochastic integrators—e.g. the second-order Heun scheme of EDM [23]—can replace Euler–Maruyama as decribed in the next subsection C.3 below.
@@
Focus on a single physical snapshot ${\bm{z}}$ as it is refined across windows. Let $k$ be the rollout step at which ${\bm{z}}$ is finalised. The snapshot first appears at rollout step $k-W+1$ (relative position $w=W$) and moves to $w=1$ at step $k$. Define $s_{w}:=\bar{\sigma}_{w}^{\text{}}(0)$ as the initial noise level when the snapshot is at position $w$, and treat $w$ as a continuous variable on the normalised interval $[0,1]$ (with $w\!\searrow\!0$ as refinement proceeds). Let ${\bm{z}}_{w}$ denote the state of the tracked snapshot when its current noise level is $s_{w}$. Its evolution is governed by

### v3

@@ Results.
An evaluation of ERDM’s performance, illustrated in Fig.˜5, reveals its consistent superiority in relative CRPS skill over our primary EDM baseline, by up to $10\%$. Similarly, ERDM demonstrates a clear advantage over related methods, significantly outperforming Graph-EFM on all four CRPS metrics reported in their study [oskarsson2024probabilistic]. Against more computationally demanding models such as IFS ENS and NeuralGCM, our approach is broadly competitive. Its primary limitation is a noted weakness in some short-range forecasts compared to IFS ENS, which we attribute to our EDM-based initialization strategy and a backbone architecture not fully specialized for weather prediction. In terms of spread-skill ratio, ERDM, along with IFS ENS, delivers the most calibrated forecasts (see Fig.˜12). In contrast, EDM and NeuralGCM often produce underdispersed short-range forecasts. A crucial aspect of evaluating any ensemble forecast is its physical realism. In this regard, ERDM performs exceptionally well. As depicted in Fig.˜6, its normalized power spectra are on par with those of the physics-based IFS ENS. This is a significant achievement, as many ML models typically struggle to reproduce such physically consistent spectra—a challenge evident with NeuralGCM in Fig.˜6 and documented by rasp2023weatherbench2. See Appendix˜F for extended results, including analyses of extra variables, spread-skill ratios, visualizations, and an EDM vs. ERDM score card.
@@ 5.4 Ablations
We conducted a comprehensive ablation study on the Navier-Stokes task to validate ERDM’s design. Our findings reveal that several architectural and training strategies are fundamental to its success; their removal causes performance to collapse, often falling below that of the single-step EDM baseline. In contrast, the model is robust to certain hyperparameters, showing little sensitivity to the window size, $W$, or the noise bounds, $\sigma_{\text{min}}$ and $\sigma_{\text{max}}$, within a reasonable range. We focus on the most impactful ablations below and provide a full analysis in Appendix˜E.
@@
    Appropriate progressive noise schedule & fixed training: The design of the noise schedule is paramount. Default EDM hyperparameters ($\rho=7$) prove suboptimal for ERDM, yielding $2\times$ worse CRPS than our proposed default schedule ($\rho=-10$). Performance is robust within a reasonable range of $\rho\in[-30,-5]$, see Fig.˜2 for intuition on why. Furthermore, training with a fixed, optimized noise schedule is critical; randomizing the schedule during training degrades performance by nearly $2\times$. Despite this, such randomized training can serve as a heuristic to identify an effective fixed schedule for subsequent retraining. This finding contrasts with the randomized training procedure from chen2024diffusionforcing.
@@
    Strategic loss weighting: Reweighing the losses based on the lognormal probability density function of noise levels, $f(\sigma)$, as proposed, is crucial. Removing this weighting causes a $2\times$ performance drop. We generally found that a larger mean of the lognormal distribution, $P_{\text{mean}}>0$, than EDM’s default, 1.2-1.2, was necessary for ERDM to achieve optimal results, confirming the importance of tuning here.
@@
    Dedicated spatiotemporal architecture: A crucial element is the use of a proper architecture that explicitly models spatiotemporal dependencies. Naively stacking the time dimension into the channel dimension of a 2D architecture results in severe $4\times$ performance degradation, confirming the need for our bespoke temporal architecture.
5.5 Computational complexity
To analyze the computational demands of ERDM relative to a standard autoregressive EDM, we benchmarked key efficiency metrics on a single A100 GPU. The results, summarized in Table˜1, correspond to generating a 30-step (15-day), 5-member ensemble weather forecast. The primary trade-off lies in the architectural design. ERDM’s hybrid 3D denoiser is inherently more memory-intensive than the 2D architecture used by EDM, requiring more than twice as much GPU memory. However, the rolling window mechanism makes ERDM significantly more efficient in terms of Neural Function Evaluations (NFEs). ERDM requires $5\times$ fewer NFEs to generate the full 15-day forecast than the step-by-step EDM. Thus, despite the higher cost per step, ERDM’s total inference time (including the initialization cost) is competitive with, and even slightly faster than, EDM’s.
Table 1: Efficiency metrics for ERDM vs. EDM. All measurements were performed on a single A100 GPU using mixed precision. Inference figures correspond to a 15-day (30 time steps), 5-member ensemble weather forecast with second-order Heun solver. Training uses a batch size of 1. NFEs are Neural Function Evaluations.
GPU Memory (GB)
Model Inference NFEs Inference Time (s) Inference Training
EDM 600 237 21 19
ERDM 120 209 49 53
@@ 6 Conclusion
We introduced the Elucidated Rolling Diffusion Model (ERDM), a diffusion framework for long-range probabilistic forecasting in complex scientific systems. ERDM adapts EDM diffusion to sequential data by integrating a progressive temporal noise schedule and snapshot-dependent preconditioning, enabling it to explicitly model the increasing uncertainty of chaotic dynamics. Our ablation studies validate that these components, combined with a bespoke spatiotemporal architecture, are synergistic and essential for strong performance. On challenging benchmarks like ERA5 weather data and Navier-Stokes simulations, ERDM consistently outperforms relevant baselines in probabilistic metrics such as CRPS, calibration, and physical realism–particularly over extended forecast horizons. Future work could improve computational efficiency through latent-space modeling or explore the framework’s applicability to a wider range of physical systems.
Limitations.
The primary limitation of ERDM is the computational cost of its 3D denoiser architecture. While ERDM requires fewer sampling steps than autoregressive baselines (e.g., $40\%$ less on ERA5), each step is more expensive, resulting in a comparable total inference time but significantly higher memory usage. This memory complexity poses a key barrier to scaling to higher resolutions, though techniques like gradient checkpointing or latent diffusion could offer a path forward. Beyond computational hurdles, ERDM’s short-range weather forecasts are not yet on par with leading operational models like IFS ENS. The framework is also constrained by its reliance on an external model for initialization and its use of an explicit noise-level loss weighting, which may be suboptimal compared to importance sampling [dieleman2024schedules], offering exciting directions for future work.
@@ Acknowledgements
S.R.C. acknowledges generous support from the summer internship and collaboration at NVIDIA. This research used resources from the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility, using NERSC awards DDR-ERCAP0034142, ASCR-ERCAP0033209, and EESSD-ERCAP0033799. This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, and NSF Grants #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069, DARPA AIE FoundSci and DARPA YFA. We are grateful to the anonymous reviewers for their valuable feedback that helped strengthen this work.
@@ Appendix B Background on Diffusion Models
Conversely, the backward diffusion process operates by initially sampling noise, represented as ${{\bm{x}}}_{0}$, from a prior distribution, typically $\mathcal{N}(\mathbf{0},\mathbf{G}(0))$, where $\mathbf{G}(0)$ corresponds to the maximum noise level. The process then focuses on denoising this sample through a sequence, ${{\bm{x}}}_{i}$, characterized by a sequence of decreasing noise covariance matrices: $\mathbf{G}(t_{0})\succ\mathbf{G}(t_{1})\succ\ldots\succ\mathbf{G}(t_{N})\approx\mathbf{0}$ (in the Loewner order), where $t_{0}=0$. Each intermediate sample ${{\bm{x}}}_{i}$ is drawn from ${{\bm{x}}}_{i}\sim p({{\bm{x}}}_{i};\mathbf{G}(t_{i}))$. The terminal sample of the backward process, ${{\bm{x}}}_{N}$, is expected to approximate the original data distribution $p_{{\rm data}}({{\bm{y}}})$. Note that for the sake of consistency with our ERDM formulation, we reversed the diffusion time order. That is, $t=0$ (11) corresponds to maximal (minimal) noise levels, which is sometimes described in opposite order in the diffusion literature.
SDE formulation. To present the forward and backward processes rigorously, they can be captured via stochastic differential equations (SDEs). Such SDEs ensure that the sample, ${\bm{x}}(t)$, aligns with the designated data distribution, $p({\bm{x}};\mathbf{G}(t))$, as it evolves through diffusion time $t$ [song2021scoreSDE]. The Elucidated Diffusion Model (EDM) by karras2022edm provides a principled design for diffusion models, typically based on a scalar noise schedule $\sigma(t)$ (leading to isotropic noise $\sigma^{2}(t)\mathbf{I}$). The formulation presented here extends these principles to accommodate a general, potentially anisotropic, time-varying noise covariance matrix $\mathbf{G}(t)$. Let $\dot{\mathbf{G}}(t)=d\mathbf{G}(t)/dt$ be the time derivative of the noise covariance matrix. We assume $\dot{\mathbf{G}}(t)$ is positive semi-definite for all $t$, and let $\mathbf{K}(t)$ be any matrix such that $\mathbf{K}(t)\mathbf{K}(t)^{T}=\dot{\mathbf{G}}(t)$ (e.g., via Cholesky decomposition if $\dot{\mathbf{G}}(t)$ is strictly positive definite, or other matrix square roots). The forward SDE, describing the process of adding noise such that ${\bm{x}}(t)-{\bm{x}}(1)\sim\mathcal{N}(\mathbf{0},\mathbf{G}(t))$ (assuming $\mathbf{G}(1)=\mathbf{0},$ and ${\bm{x}}(1)\sim p_{\text{data}}({\bm{y}})$), is:
@@
where $\boldsymbol{\omega}(t)$ is a standard $d$-dimensional Wiener process. The corresponding backward (reverse-time) SDE is given by [song2021scoreSDE]:
@@
where $\Delta\mathbf{G}(\tau_{i})=\mathbf{G}(\tau_{i})-\mathbf{G}(\tau_{i+1})>\mathbf{0}$, we have used the approximation $\dot{\mathbf{G}}(\tau_{i})\Delta\tau_{i}\approx\mathbf{G}(\tau_{i})-\mathbf{G}(\tau_{i+1})$ (assuming $\Delta\mathbf{G}(\tau_{i})$ is small), and $\mathbf{z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a standard Gaussian noise sample. The term $\hat{\nabla}_{{{\bm{x}}}}\log p({\bm{x}}_{\tau_{i}};\mathbf{G}(\tau_{i}))$ is the score estimated using the trained denoiser:
$\hat{\nabla}_{{{\bm{x}}}}\log p({\bm{x}}_{\tau_{i}};\mathbf{G}(\tau_{i}))=(\mathbf{G}(\tau_{i}))^{-1}(D_{\theta}({\bm{x}}_{\tau_{i}};\mathbf{G}(\tau_{i}))-{\bm{x}}_{\tau_{i}}).$ (10)
More sophisticated samplers, such as those proposed by karras2022edm (e.g., second-order Heun or DPM-Solver++), can also be adapted for the general covariance case to improve sample quality or reduce the number of sampling steps $N$. These often involve more complex update rules, potentially incorporating predictor-corrector steps or higher-order approximations of the SDE.
@@ C.1 Probability flow ODE derivation
Recall that ERDM operated on noisy sequences $\bar{\bm{x}}_{1:W}$ of size $W$, with each $\bar{\bm{x}}_{w}\in\mathbb{R}^{D}$ corrupted according to increasing noise level standard deviations $0\approx\bar{\sigma}_{1}(t)<\bar{\sigma}_{2}(t)<\dots<\bar{\sigma}_{W}(t)$. The noise covariance matrix for the entire window $\bar{\bm{x}}_{1:W}$ at diffusion time t{t} is block-diagonal:
@@
where $\boldsymbol{\omega}({t})$ is a standard $WD$-dimensional Wiener process. The corresponding backward (reverse-time) SDE, which is the primary SDE used for generation, is [song2021scoreSDE]:
@@
where $D_{\theta}(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))_{w}$ is the $w$-th snapshot output of the denoiser $D_{\theta}$, which is conditioned on the full noisy window $\bar{\bm{x}}$ and the vector of current noise standard deviations $\bm{\bar{\sigma}}({t})$. The probability flow ODE, which provides a deterministic path for generation, is obtained by removing the stochastic term from the backward SDE and adjusting the drift (see song2021scoreSDE, Eq. (14) and karras2022edm, Eq. (5)):
@@
Using the specific EDM parameterization where the ODE drift is $\mathbf{D}({t})\nabla_{\bar{\bm{x}}}\log p(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))$ with $\mathbf{D}({t})=\text{diag}(\bar{\sigma}_{1}({t})\dot{\sigma}_{1}({t})\mathbf{I}_{D},\ldots,\bar{\sigma}_{W}({t})\dot{\sigma}_{W}({t})\mathbf{I}_{D})$, the ODE becomes:
@@ C.2 Sampling
ERDM generates sequences with a sliding window of size $W$. Generating each window requires solving the backward SDE in equation 17 (or ODE in equation 20). In each iteration $k$ where the SDE is fully solved, ERDM operates on the noisy window $\bar{\bm{x}}^{(k)}_{1:W}$. After solving the SDE backwards $k\geq 1$ times, the first snapshot of the window, $\bar{\bm{x}}^{(k)}_{1}$ (with relative window index $w=1$), corresponds to a forecast of ${\bm{y}}_{k}$, where $k$ is the absolute time step of the forecast time dimension. Solving the SDE requires discretizing it into $N_{k}$ diffusion “time” steps, each of size $\Delta{t}_{k}:=\frac{1}{N_{k}}$. In general, $N_{k}$ may vary across iterations. To simplify our illustrations of the sampling process in the following, we assume that $N_{k}$ and $\Delta{t}_{k}$ are identical for all $k$, denoted as $N$ and $\Delta{t}$ respectively. This can be achieved by setting $N$ in our sampling algorithms to a positive integer. Note that in practice $N$ need not be an integer (e.g., we use $N=1.25$ for our Navier-Stokes experiments), which is accounted for in our sampling algorithms. With this simplification, the SDE evolves, for all $k$, following $\bm{\bar{\sigma}}(t_{1})\rightarrow\bm{\bar{\sigma}}(t_{2})\rightarrow\dots\rightarrow\bm{\bar{\sigma}}(t_{N})$, where $t_{i+1}=t_{i}+\Delta{t}$ for $i\geq 1$, $t_{1}=0$, and $t_{N}=1$. For example, if $N=1$ the SDE is solved from $\bm{\bar{\sigma}}(0)$ to $\bm{\bar{\sigma}}(1)$ in one sampling step, while if $N=2$ it would be solved following $\bm{\bar{\sigma}}(0)\rightarrow\bm{\bar{\sigma}}(\frac{1}{2})\rightarrow\bm{\bar{\sigma}}(1)$ in two sampling steps (see illustration in Fig.˜1). At the end of the $N$ sampling steps, $\bar{\sigma}_{1}(1)=\sigma_{\text{min}}$ and the first snapshot in the window can be emitted as the forecast for lead time $k$, $\bar{\bm{x}}^{(k)}_{1}\approx{\bm{y}}_{k}$. In the next iteration of $N$ sampling steps, the next snapshot $k+1$ will be predicted by operating on a new window, $\bar{\bm{x}}^{(k+1)}_{1:W}$. Because $\bar{\sigma}_{w+1}(1)=\bar{\sigma}_{w}(0)$ for $w\geq 1$, this new window can be constructed by shifting the noisy parts of the old window by one position. That is, $\bar{\bm{x}}^{(k+1)}_{w}:=\bar{\bm{x}}^{(k)}_{w+1}$ for $w<W$, and a new pure-noise snapshot is appended as $\bar{\bm{x}}^{(k+1)}_{W}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I})$. The new window is thus noised according to standard deviations $\bm{\bar{\sigma}}(0)$ and the process can be restarted and repeated indefinitely, as illustrated in Fig.˜1.
@@
The ODE is thus solved in parallel for all snapshots in the window. When $\bar{\bm{x}}(N)$ is reached, its noise levels correspond to $\bm{\bar{\sigma}}(N\Delta{t})=\bm{\bar{\sigma}}(1)$. Thus, the first snapshot is emitted, and the window is shifted as describe above. Higher-order stochastic integrators—e.g. the second-order Heun scheme of EDM [karras2022edm]—can replace Euler–Maruyama as decribed in the next subsection C.3 below.
@@
Focus on a single physical snapshot $\bm{z}$} as it is refined across windows. Let $k$ be the rollout step at which $\bm{z}$} is finalised. The snapshot first appears at rollout step kW+1k-W+1 (relative position $w=W$) and moves to $w=1$ at step $k$. Define $s_{w}:=\bar{\sigma}_{w}^{\text{}}(0)$ as the initial noise level when the snapshot is at position $w$, and treat $w$ as a continuous variable on the normalised interval $[0,1]$ (with $w\!\searrow\!0$ as refinement proceeds). Let ${\bm{z}}_{w}$ denote the state of the tracked snapshot when its current noise level is $s_{w}$. Its evolution is governed by